# Learning Operators through Coefficient Mappings in Fixed Basis Spaces

### v1

@@ Abstract
Operator learning has emerged as a powerful paradigm for approximating solution operators of partial differential equations (PDEs) and other functional mappings. Classical approaches typically adopt a pointwise-to-pointwise framework, where input functions are sampled at prescribed locations and mapped directly to solution values. We propose the Fixed-Basis Coefficient to Coefficient Operator Network (FB-C2CNet), which learns operators in the coefficient space induced by prescribed basis functions. In this framework, the input function is projected onto a fixed set of basis functions (e.g., random features or finite element bases), and the neural operator predicts the coefficients of the solution function in the same or another basis. By decoupling basis selection from network training, FB-C2CNet reduces training complexity, enables systematic analysis of how basis choice affects approximation accuracy, and clarifies what properties of coefficient spaces (such as effective rank and coefficient variations) are critical for generalization. Numerical experiments on Darcy flow, Poisson equations in regular, complex, and high-dimensional domains, and elasticity problems demonstrate that FB-C2CNet achieves high accuracy and computational efficiency, showing its strong potential for practical operator learning tasks.
@@ 1 Introduction
Deep learning has recently emerged as a powerful paradigm in scientific computing, with applications ranging from solving differential equations to accelerating large-scale simulations and data-driven modeling [1, 2, 3, 4]. Among various methodologies, operator learning has garnered significant attention, as it provides a principled framework for approximating mappings between infinite-dimensional function spaces directly. Unlike neural PDE solvers such as PINNs [5], DeepRitz [6], and WANs [7], which approximate the solution of a partial differential equation (PDE) at finitely many spatial or temporal points, operator learning  [8, 9, 10] aims to approximate the entire mapping from input functions—such as initial conditions, boundary conditions, or source terms—to output functions, typically the corresponding PDE solutions. For instance, consider the operator $G:(a,f)\mapsto u,$ which maps a coefficient function $a$ and a source term $f$ to the solution $u$ of the elliptic PDE $\nabla\cdot(a\nabla u)=f,$ subject to appropriate boundary conditions. This example will serve as a recurring case study throughout this paper. This operator-based perspective represents a paradigm shift: once trained, operator learning models can serve as fast and generalizable solvers applicable to a wide range of problem settings. This stands in contrast to conventional numerical algorithms, which typically rely on case-specific discretizations and require repeated, computationally intensive simulations for each new instance.
@@
Despite its great potential, operator learning faces several fundamental challenges. One of the most critical issues is the discretization of infinite-dimensional function spaces. In other words, a key question is: what should serve as the input and output representations of the neural network? To make the problem computationally tractable, these infinite-dimensional spaces must be approximated by finite-dimensional subspaces. In most existing operator learning frameworks [8, 9, 10, 11, 12, 13], functions are discretized by their nodal values at a collection of collocation points, which are then used as the neural network input and output. While this pointwise encoding is straightforward and widely adopted, it often suffers from drawbacks such as mesh dependence and deteriorating generalization as the resolution increases.
@@
In this work, we propose a fixed basis coefficient-to-coefficient (FB-C2C) operator learning framework. The key idea is to represent both input and output functions in terms of carefully chosen basis functions, and to train a neural network to learn the mapping between their expansion coefficients. Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be an operator that maps an input function $f\in\mathcal{X}$ to an output function $G(f)\in\mathcal{Y}$. We approximate the input function space $\mathcal{X}$ by a finite-dimensional subspace $V_{\mathrm{in}}=\text{span}\{\phi_{j}(x)\}_{j=1}^{m_{1}}$ and the output function space $\mathcal{Y}$ by $V_{\mathrm{out}}=\text{span}\{\psi_{j}(y)\}_{j=1}^{m_{2}}$. For each input function $f(x)$, we approximate it as
@@ 1.1 Related work
In recent years, leveraging neural network architectures for solving partial differential equations (PDEs) has attracted substantial attention in the scientific computing community. Such approaches are particularly promising in handling problems defined on complex domains and in incorporating empirical data into mathematical models. From a theoretical standpoint, neural networks have the capacity to overcome the curse of dimensionality in high-dimensional PDE problems [14, 15, 16, 17, 18]. Broadly speaking, existing neural network–based methods for solving PDEs can be categorized into two major classes.
The first is PDE solution approximation, where neural networks are trained to approximate the solution of a PDE directly. Representative approaches in this category include Physics-Informed Neural Networks (PINNs) [5, 19, 20], the Deep Ritz Method [6], Weak Adversarial Networks (WANs)  [7, 21], and random feature models [22, 23, 24, 13]. These methods typically design specialized loss functions that encode the underlying PDE constraints, boundary conditions, or variational formulations, thereby guiding the neural network toward solutions consistent with the governing equations. Such approaches have demonstrated notable success in forward problems, inverse problems, and data assimilation tasks. However, their efficiency often deteriorates as the dimensionality of the problem grows, and training may become prohibitively expensive or unstable for highly nonlinear systems.
The second class is operator learning, which shifts the perspective from approximating a single PDE solution to learning the underlying solution operator. In this framework, the goal is to approximate mappings from input functions—such as coefficients, source terms, or boundary conditions—to corresponding solution functions. This operator-centric viewpoint not only enables rapid solution of new problem instances once the operator is learned, but also facilitates generalization across families of PDEs. Representative methods include DeepONet [8] and the Fourier Neural Operator (FNO) [9], which introduced new architectures for handling function-to-function mappings efficiently. Building upon these seminal works, a growing number of operator-learning–based approaches have been developed to approximate increasingly complex classes of operators, ranging from multi-scale operators [25], to domain decomposition operators [26], to operators arising in physical and material sciences [27, 28]. These developments underscore the growing importance of operator learning as a paradigm that bridges classical numerical analysis with modern machine learning. Our work is centered on operator learning, and in what follows, we review existing literature and approaches that are closely connected to the present study.
@@
DeepONet [8] is one of the earliest and most influential works on operator learning. Building on the Universal Approximation Theorem for Operators, it introduces the framework:
@@
where the Branch Net takes as input the function values at sensor points $f(x_{j})$, while the Trunk Net takes spatial coordinates and provides basis functions. The final output is obtained by combining the coefficients from the Branch Net with the basis from the Trunk Net. Following this, POD-DeepONet [10] replaced the Trunk Net with Proper Orthogonal Decomposition (POD) modes extracted from training data, while keeping the Branch Net to learn their coefficients. Later studies further explored alternative trunk representations, including fixed bases from Extreme Learning Machines (ELM) [11], finite element shape functions in Mesh-Informed Neural Networks (MINNs) [12], and random feature models with partition-of-unity strategies [13]. Despite these variations, the Branch Net still relies on $f(x_{j})$ as input. However, this design suffers from scalability issues: in higher dimensions, the number of sensor points becomes large, leading to high-dimensional inputs, larger networks, slower training, and potentially weaker generalization.
To address the challenges posed by high-dimensional inputs in operator learning, recent advances have explored integrating neural networks with principled dimensionality reduction strategies [29, 30, 31, 32, 33, 34]. Among these, three works are particularly relevant to our approach. The Basis Operator Network [35] represents input functions using basis coefficients: the Branch Net takes as input the coefficients of the selected basis functions, while the Trunk Net encodes spatial locations, with orthogonality of the Trunk Net preserved throughout training. A related idea is pursued in Basis-to-Basis Operator Learning Using Function Encoders [36], where the Trunk and Branch Nets are trained separately. In this framework, the trained Trunk Net is treated as a function encoder, and its output coefficients are subsequently used as inputs to the Branch Net. In the Resolution-Independent Neural Operator paper [37], the trunk network selects basis functions from a predefined dictionary of functions, and two dictionary learning algorithms are employed to adaptively learn a set of suitable continuous basis functions. The main limitation of these approaches lies in the fact that their basis functions must be obtained through training, which is both time-consuming and prone to instability. Our proposed approach is to directly select a function space—such as finite element basis functions or random feature models—as the function encoder, and then compute the coefficients of the input function with respect to these basis functions using the least squares method. This makes our method more effective, leading to shorter overall training time and improved stability. The coefficients are subsequently used as the input to the neural operator.
@@
    We investigate different choices of approximation function spaces, including random feature models and finite element bases, and further discuss strategies for computing the coefficients in a manner that facilitates learning by the neural operator. In particular, we analyze which characteristics of the coefficient representations are most critical for generalization, highlighting the role of effective rank and coefficient variations in producing representations that are both expressive and stable.
@@
    We demonstrate the effectiveness of the proposed approach through extensive numerical experiments, including Darcy flow in 1D/2D regular domains, Poisson equations in complex domains, the Elastic equation, Darcy flow in complex domains, and high-dimensional Poisson equations, showing that our method achieves both high efficiency and accuracy.
@@
    We further generalize the framework to multi-input/output operator learning, involving mappings between multiple input and output functions. We compare scalar-valued representations, obtained by concatenating input/output coefficients, with vector-valued constructions that employ shared vector-valued basis functions defined over all components. Numerical experiments show that the RFM-based vector-valued formulation consistently achieves higher stability and accuracy.
The remainder of the paper is organized as follows. Section 2 provides a detailed introduction to the proposed fixed basis coefficient-to-coefficient operator learning framework. Section 3 discusses the design choices of key components, including the selection of basis function spaces and strategies for coefficient computation. Section 4 presents numerical experiments, covering Darcy flow, Poisson equations on both regular and complex domains, as well as high-dimensional settings, and elastic plate equation to demonstrate the effectiveness, efficiency, and accuracy of our method. Section 5 concludes with a discussion and outlines directions for future work.
@@ 2.2 Fixed-Basis Coefficient to Coefficient Operator Network (FB-C2CNet)
Our method consists of two main stages. The first stage is fixed basis function selection. In this stage, we specify two basis function systems, denoted as $V_{\text{in}}=\{\phi_{i}(\mathbf{x})\}_{i=1}^{m_{1}}$ and $V_{\text{out}}=\{\psi_{i}(\mathbf{y})\}_{i=1}^{m_{2}}$, which are employed to approximate the input function space $\mathcal{X}$ and the output function space $\mathcal{Y}$, respectively. In the subsequent step, a given input function $f(\mathbf{x})\in\mathcal{X}$ is represented in terms of the basis $V_{\text{in}}=\{\phi_{i}(\mathbf{x}),\phi_{2}(\mathbf{x}),\dots,\phi_{m_{1}}(\mathbf{x})\}$, i.e.,
@@
The loss function employed to train FB-C2CNet is the relative loss, defined as
$L(\theta)=\frac{1}{M}\sum_{k=1}^{M}\frac{\left\|NN_{\theta}(\mathbf{a}^{(k)})\cdot\mathbf{\Psi}-\mathbf{u}^{(k)}\right\|_{2}}{\left\|\mathbf{u}^{(k)}\right\|_{2}},$ (7)
@@
We adopt the mean relative error loss instead of the plain mean squared error (MSE) in order to normalize the approximation error with respect to the magnitude of the ground truth. This normalization ensures scale-invariance across different training samples: functions with larger amplitudes do not dominate the optimization process, while functions with smaller amplitudes are not overlooked. Such a formulation is particularly important in operator learning, where the outputs $\{\mathbf{u}^{(k)}\}_{k=1}^{M}$ may vary significantly in scale across different samples. By employing the relative loss, FB-C2CNet achieves a more balanced training process and yields a more robust generalization performance. The overall procedure is illustrated in Figure 2 and Algorithm.
@@ 3 Basis Systems and Coefficient Estimation
$NN_{\bm{\theta}}:\mathbb{R}^{m_{1}}\to\mathbb{R}^{m_{2}},$
@@ 3.1 Fixed Basis Systems
In one dimension, $\psi_{n}$ can be either the characteristic function
@@
In addition to FEM and random feature models, a variety of other basis function systems can be employed for constructing approximation spaces. Classical spectral methods, for instance, utilize global orthogonal polynomials such as Chebyshev or Legendre polynomials, which offer exponential convergence rates for smooth solutions [46, 47]. Alternatively, radial basis functions (RBFs), including Gaussian, multiquadric, and inverse multiquadric kernels, provide meshfree approximations that are particularly effective for scattered data and irregular domains [48, 49]. Wavelet bases have also been explored, offering multiresolution representations with good localization in both space and frequency [50]. Each of these choices comes with distinct advantages and trade-offs in terms of accuracy, computational cost, and adaptability to complex geometries, and they can be flexibly integrated into operator-learning frameworks depending on the problem setting.
Figure 3: *

### v2

@@ Abstract
Operator learning has emerged as a promising paradigm for approximating solution operators of partial differential equations (PDEs). However, conventional approaches typically rely on pointwise function discretizations, which often suffer from the curse of dimensionality, mesh dependence, and prohibitive training costs in high-resolution settings. To address these challenges, we propose the Fixed-Basis Coefficient to Coefficient Operator Network (FB-C2CNet), a novel framework that learns the operator mapping within the coefficient spaces induced by prescribed, fixed basis functions. Unlike existing methods that learn basis functions dynamically or rely on extensive sensor grids, FB-C2CNet encodes input functions onto a fixed set of basis functions (such as random features or finite element bases) and employs a neural network to predict the expansion coefficients of the solution. By decoupling basis selection from network training, this formulation significantly reduces the dimensionality of the input-output spaces and the number of trainable parameters. We further introduce metrics such as effective rank to analyze how the spectral properties of the coefficient space influence generalization performance. Extensive numerical experiments across a wide spectrum of benchmarks—including linear, nonlinear, and high-dimensional problems—demonstrate that FB-C2CNet achieves competitive predictive accuracy while reducing training time by orders of magnitude compared to conventional neural operators.
@@ 1 Introduction
Deep learning has recently emerged as a powerful paradigm in scientific computing, with applications ranging from solving differential equations to accelerating large-scale simulations and data-driven modeling [37, 5, 35, 29]. Among various methodologies, operator learning has garnered significant attention, as it provides a principled framework for approximating mappings between infinite-dimensional function spaces directly. Unlike neural PDE solvers such as Physics-Informed Neural Networks (PINNs) [38, 27, 13], the Deep Ritz Method (DeepRitz) [16], Weak Adversarial Networks (WANs)  [45, 3], and random feature models (RFM) [10, 15, 41, 9], which approximate the solution of a partial differential equation (PDE) at finitely many spatial or temporal points, operator learning [33, 31, 34] aims to approximate the entire mapping from input functions such as initial conditions, boundary conditions, or source terms to output functions, typically the corresponding PDE solutions. For instance, consider the operator $G:(a,f)\mapsto u,$ which maps a coefficient function $a$ and a source term $f$ to the solution $u$ of the elliptic PDE: $\nabla\cdot(a\nabla u)=f,$ subject to appropriate boundary conditions. This example will be used as a running illustrative case throughout the paper. This operator-based perspective represents a paradigm shift: once trained, operator learning models can serve as fast and generalizable solvers applicable to a wide range of problem settings. This stands in contrast to conventional numerical algorithms, which typically rely on case-specific discretizations and require repeated, computationally intensive simulations for each new instance.
@@
Despite its great potential, operator learning faces several fundamental challenges. One of the most critical issues lies in the discretization of infinite-dimensional function spaces. In other words, a key question is: what should serve as the input and output representations of a neural operator? To make the problem computationally tractable, these infinite-dimensional spaces are typically approximated by finite-dimensional subspaces. In most existing operator learning frameworks [33, 31, 34, 40, 44, 9], functions are discretized through their nodal values at a collection of collocation points, which are then used as the neural network input and output. While this pointwise encoding is straightforward and widely adopted, it suffers from several inherent drawbacks, including strong mesh dependence and degraded generalization performance as the spatial resolution increases. Moreover, in two- and three-dimensional problems, achieving adequate accuracy often requires a large number of grid points, leading to extremely high-dimensional inputs. This, in turn, results in neural networks with a prohibitively large number of parameters, significantly increased training cost, and long training times.
@@
In this work, we propose a fixed basis coefficient-to-coefficient (FB-C2C) operator learning framework. The key idea is to represent both input and output functions in terms of chosen basis functions, and to train a neural network to learn the mapping between their expansion coefficients. Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be an operator that maps an input function $f\in\mathcal{X}$ to an output function $G(f)\in\mathcal{Y}$. We approximate the input function space $\mathcal{X}$ by a finite-dimensional subspace $V_{\mathrm{in}}=\text{span}\{\phi_{j}(x)\}_{j=1}^{m_{1}}$ and the output function space $\mathcal{Y}$ by $V_{\mathrm{out}}=\text{span}\{\psi_{j}(y)\}_{j=1}^{m_{2}}$. For each input function $f(x)$, we approximate it as
@@ 1.1 Related work
Neural network–based methods for solving partial differential equations (PDEs) have been extensively studied in recent years. Existing approaches can be broadly categorized into two classes.
The first class focuses on direct approximation of PDE solutions, where neural networks are trained to represent individual solutions. Representative methods include PINNs [38, 27], the Deep Ritz method [16], WANs [45, 3], and RFM [10, 15]. These methods typically incorporate PDE constraints through loss functions derived from strong, weak, or variational formulations.
The second class is operator learning, which aims to approximate solution operators mapping input functions—such as coefficients, source terms, or boundary conditions—to solution functions. Representative approaches include DeepONet [33] and the Fourier Neural Operator (FNO) [31]. A variety of extensions have since been proposed to address multiscale problems, domain decomposition, and complex physical systems [19, 28, 30]. The present work is situated within the operator learning framework.
@@
DeepONet [33] is one of the earliest and most influential works on operator learning. Building on the Universal Approximation Theorem [22] for Operators, it introduces the framework:
@@
where the Branch Net takes as input the function values at sensor points $f(x_{j})$, while the Trunk Net takes spatial coordinates and outputs basis functions. In this paper, we refer to such approaches as point-to-coefficient (P2C) methods, emphasizing that the Branch Net maps pointwise function evaluations to coefficient representations. The final output is obtained by combining the coefficients from the Branch Net with the basis from the Trunk Net. Following this, POD-DeepONet [34] replaced the Trunk Net with Proper Orthogonal Decomposition (POD) modes extracted from training data, while keeping the Branch Net to learn their coefficients. Later studies further explored alternative trunk representations, including fixed bases from Extreme Learning Machines (ELM) [40], finite element shape functions in Mesh-Informed Neural Networks (MINNs) [44], and random feature models with partition-of-unity strategies [9]. Despite these variations, the Branch Net still relies on $f(x_{j})$ as input. However, this design suffers from scalability issues: in higher dimensions, the number of sensor points becomes large, leading to high-dimensional inputs, larger networks, slower training, and potentially weaker generalization.
To address the challenges posed by high-dimensional inputs in operator learning, recent work has explored integrating neural networks with principled dimensionality reduction strategies [1, 4, 42, 20, 21, 46]. Among these, three approaches are particularly relevant to our work. The Basis Operator Network (BasisONet) [24] represents input functions using coefficients with respect to a prescribed basis. In this framework, the Branch Net takes the basis coefficients as input, while the Trunk Net encodes spatial locations, and both networks are trained jointly, with orthogonality constraints imposed on the Trunk Net throughout training. Another closely related approach is the Basis-to-Basis (B2B) operator learning method [25], in which the Trunk and Branch Nets are trained separately. In the Resolution-Independent Neural Operator framework [2], the trunk network selects basis functions from a predefined dictionary, and dictionary learning techniques are employed to adaptively construct suitable continuous basis functions.
A common limitation of these approaches is that the underlying basis functions are learned during training, which can be computationally expensive and potentially unstable. In contrast, our approach directly selects a predefined function space—such as finite element bases or random feature models—as the function encoder, and computes the corresponding expansion coefficients via a least-squares projection, which we refer to as the encoding step. These coefficients are then used as inputs to the neural operator, leading to improved training stability and reduced overall computational cost.
@@
    We investigate different choices of approximation function spaces, including random feature models and finite element bases, and study how the resulting coefficient representations influence the generalization performance of neural operators. In particular, we introduce effective rank and coefficient variation to quantify whether the distribution of the input coefficients is conducive to learning representations that enable robust generalization.
@@
    We validate the effectiveness of the proposed FB-C2CNet framework through extensive numerical experiments across a diverse set of benchmark problems spanning multiple PDE classes, including steady-state elliptic equations, time-dependent nonlinear PDEs, Helmholtz equations with only weak solutions, and inverse boundary problems. Across all benchmarks, FB-C2CNet consistently achieves competitive or superior accuracy compared to existing neural operator methods, while requiring substantially reduced training cost.
@@
    We further extend the proposed framework to multi-input/output operator learning, where mappings are defined between multiple input and output functions. We compare scalar-valued coefficient representations with vector-valued constructions based on shared vector-valued basis functions, and demonstrate that the RFM-based vector-valued formulation consistently achieves improved stability and accuracy. Moreover, an inverse boundary problem for the Stokes system is used as a representative application, illustrating the effectiveness of the proposed framework in realistic inverse problem settings.
The remainder of the paper is organized as follows. Section 2 provides a detailed introduction to the proposed fixed basis coefficient-to-coefficient operator learning framework. Section 3 discusses the design choices of key components, including the selection of basis function spaces and strategies for coefficient computation. Section 4 presents numerical experiments demonstrating the effectiveness, efficiency, and accuracy of the proposed method, followed by conclusions and future directions in Section 5.
@@ 2.2 Fixed-Basis Coefficient to Coefficient Operator Network (FB-C2CNet)
Our method consists of two main stages. The first stage is fixed basis function selection. In this stage, we specify two basis function systems, denoted as $V_{\text{in}}=\text{span}\{\phi_{i}(\mathbf{x})\}_{i=1}^{m_{1}}$ and $V_{\text{out}}=\text{span}\{\psi_{i}(\mathbf{y})\}_{i=1}^{m_{2}}$, which are employed to approximate the input function space $\mathcal{X}$ and the output function space $\mathcal{Y}$, respectively. In the subsequent step, a given input function $f(\mathbf{x})\in\mathcal{X}$ is represented in terms of the basis $V_{\text{in}}=\text{span}\{\phi_{i}(\mathbf{x}),\phi_{2}(\mathbf{x}),\dots,\phi_{m_{1}}(\mathbf{x})\}$, i.e.,
@@
The loss function employed to train FB-C2CNet is the mean relative $L^{2}$ error, defined as
$L(\bm{\theta})=\frac{1}{M}\sum_{k=1}^{M}\frac{\left\|NN_{\theta}(\mathbf{a}^{(k)})\cdot\mathbf{\Psi}-\mathbf{u}^{(k)}\right\|_{2}}{\left\|\mathbf{u}^{(k)}\right\|_{2}},$ (7)
@@
We adopt the mean relative error loss instead of the plain mean squared error (MSE) in order to normalize the approximation error with respect to the magnitude of the ground truth. This normalization ensures scale-invariance across different training samples: functions with larger amplitudes do not dominate the optimization process, while functions with smaller amplitudes are not overlooked. Such a formulation is particularly important in operator learning, where the outputs $\{\mathbf{u}^{(k)}\}_{k=1}^{M}$ may vary significantly in scale across different samples. By employing the relative loss, FB-C2CNet achieves a more balanced training process and yields a more robust generalization performance. The overall procedure is illustrated in Figure 2 and Algorithm. For clarity, the main notation used throughout the paper is summarized in Table 1.
@@
Table 1: Summary of notation and key quantities used in the paper.
Category Description Notation
Dataset Number of training samples $N_{\mathrm{train}}$
Number of testing samples $N_{\mathrm{test}}$
Number of input grid points $n_{1}$
Number of output grid points $n_{2}$
Basis representation Number of input basis functions $m_{1}$
Number of output basis functions $m_{2}$
Random Feature Model (RFM) Initial variance of random features $R_{m}$
Number of partitions of unity (PoU) $M_{p}$
Spectral cut-off parameter $\mathrm{cut}$
@@ 3 Basis Systems and Coefficient Estimation
$NN_{\bm{\theta}}:\mathbb{R}^{m_{1}}\mapsto\mathbb{R}^{m_{2}},$
@@ 3.1 Fixed Basis Systems
In one dimension, $\omega_{n}$ can be either the characteristic function
@@
In addition to FEM and random feature models, a variety of alternative basis systems can be used to construct approximation spaces. Classical spectral methods employ global orthogonal polynomials, such as Chebyshev or Legendre bases, and achieve exponential convergence for smooth solutions [6, 43]. Radial basis functions (RBFs), including Gaussian and multiquadric kernels, provide meshfree approximations well suited for scattered data and irregular domains [17, 8]. Wavelet bases offer multiresolution representations with simultaneous localization in space and frequency [14]. Each of these basis choices entails different trade-offs in accuracy, computational cost, and geometric flexibility, and can be incorporated into operator-learning frameworks as dictated by the problem setting.